Use two equations in two variables to solve the application. See Example 1. (Objective 1)
The salaries of the President and Vice President of a certain country total $682,400 a year. If the President makes $217,600 more than the Vice President, find each of their salaries.

Answer by ikleyn(52780)

.
Let x be the salary of the president and y be the salary of the Vice President.


Then from the condition, you have these two equations in two unknowns


    x + y = 682400   dollars     (1)

    x - y = 217600   dollars     (2)


Add equations (1) and (2).  You will get then


    2x = 682400 + 217600 = 900000

     x                   = 900000/2 = 450000.


Next, from equation (1), you find

    y = 682400 - 450000 = 232400.


ANSWER.  $450000 is the salary of the President  and  $232400  is the salary of the Vice President.

Solved.

---------------

If you want to see one million other similar solved problems, look into the lesson
    - Word problems that lead to a simple system of two equations in two unknowns
in this site.